### v3

@@ 4.4.3 Why is cross-event transfer challenging?
To better understand why cross-event generalization is so difficult, we explore two fundamental factors rooted in the nature of real-world disaster response.
@@
    Inconsistent damage signatures across events. Figure 9 presents violin plots of SAR backscatter values for intact, damaged, and destroyed buildings across multiple event pairs. These plots reveal two key observations. Firstly, even within the same disaster category (the first row), the pixel intensity distributions for damaged and destroyed buildings differ significantly between events. This indicates that SAR-based damage signatures are not consistent across locations, possibly due to differences in urban layout, building materials, or sensor incidence angles. Secondly, the distribution shift becomes more pronounced across different disaster types (the second row). For instance, the signatures of destroyed building in wildfires are notably different from those in volcano eruptions, which again differ from floods or hurricanes. These variations reflect fundamental differences in damage mechanisms: buildings burned in wildfires, submerged in floods, or collapsed in earthquakes leave very different patterns in SAR backscatter. Such distributional discrepancies make it extremely difficult for models to generalize from one event to another. A model trained on one disaster may learn class boundaries (e.g., between “Damaged” and “Destroyed”) that do not transfer well to another disaster, especially when the visual and physical properties of damage are fundamentally different.
    Figure 10: Comparison of models’ best performance (mIoU) on test events versus the best checkpoints selected on validation sets under the cross-event transfer setting. The farther a point lies from the diagonal line, the larger the gap between the model’s selected performance and its true upper bound.
@@
    Lack of target domain supervision for model selection. Another critical challenge in cross-event setup is the absence of labeled target domain samples for model selection. In typical domain adaptation benchmarks in computer vision, a target domain validation set is available to tune hyperparameters or select the best checkpoint. However, in disaster response scenarios, no such validation data can be assumed. Figure 10 illustrates the resulting issue: the gap between the actual best mIoU on the test event and the mIoU of the checkpoint selected using a source domain validation set for several representative models under both zero-shot and one-shot settings. In the zero-shot setting, this performance gap is substantial across most models. This indicates that relying on source domain validation leads to suboptimal model selection due to poor reflection of the target distribution. The one-shot setting helps reduce this gap by enabling limited target-aware selection. However, the issue is not entirely resolved. Even with a few labeled samples, model instability and domain shift still make selection challenging.
Together, these two findings highlight that cross-event transfer is hindered not only by semantic and statistical shifts across disaster types but also by operational constraints that prevent ideal training and tuning. Future solutions should consider both aspects, i.e., developing models that are robust to distributional variance and designing selection strategies that do not depend on target supervision, such as self-validation, early stopping heuristics, or domain-agnostic performance proxies (Yang et al., 2023).
4.5 Extended evaluation of unsupervised multimodal change detection methods
Table 11: Evaluation results of representative unsupervised multimodal change detection methods. The highest values are highlighted in purple, and the second-highest results are highlighted in teal. The accuracies on the UMCD benchmark dataset are the accuracies on the four datasets presented in Figure 12. Details of methods and benchmark datasets are introduced in Appendix G.
@@ 4.5 Extended evaluation of unsupervised multimodal change detection methods
Unsupervised multimodal change detection (UMCD) plays a crucial role in post-disaster assessment by enabling rapid analysis of affected areas without requiring labeled data or prior model training (Lv et al., 2022). The current research is mainly conducted on toy datasets with limited geographic diversity and scene complexity. These datasets mainly involve general land-cover changes. It is still unknown whether these methods can work in real disaster occurrence/building damage scenarios. Bright offers a new opportunity to evaluate UMCD methods, thereby providing insights into their robustness and scalability. Six representative UMCD methods (Sun et al., 2021; Chen et al., 2022b, 2023; Han et al., 2024; Sun et al., 2024a; Liu et al., 2025) are evaluated in this paper. The specific experimental setup is described in Appendix G.
@@
    Limitations in traditional UMCD datasets. The traditional UMCD datasets consist of only a handful of image pairs, often depicting simple land-cover changes such as urban expansion or deforestation, with relatively low-resolution imagery with limited geographic diversity, as shown in Figure 12. These datasets fail to capture the complexity and variability found in real-world disaster scenarios. In contrast, Bright provides thousands of VHR multimodal image pairs across different types of disasters, significantly increasing the diversity of test cases and making it a more challenging benchmark for UMCD model evaluation.
@@
    Interference from non-disaster changes. Unlike previous UMCD benchmarks where general land-cover changes are the primary detection target, changes such as vegetation growth, water body shifts, and urban development may introduce noise and interfere with damage detection in real disaster scenarios. While prior UMCD studies treat such changes as valid detection targets, Bright requires methods to differentiate true building structural damage from irrelevant land-cover variations, posing a unique challenge to current approaches.
@@
    Problematic evaluation protocol. Current UMCD research follows a problematic evaluation protocol where models are trained, hyperparameters are tuned, and then validated on the same dataset. This approach can actually lead to overfitting, artificially making the model look very accurate but not generalizable to real-world scenarios. In a real disaster scenario, there are no labeled samples available for hyperparameter tuning, making the current practice unrealistic. Bright exposes this limitation, as models now need to be trained or tuned before being tested directly on new, unseen data, requiring truly generalizable and adaptive learning strategies.
The results highlight the challenges of applying existing UMCD models to real-world disaster scenarios. Bright reveals significant limitations in current methodologies and also presents new opportunities for future UMCD research.
4.6 Extended evaluation of unsupervised multimodal image matching methods
@@
Finally, precise image alignment is a critical prerequisite for any multimodal EO application. Bright offers a unique opportunity to evaluate the performance of existing UMIM algorithms under realistic, large-scale disaster conditions. Due to the lack of real pixel-level ground truth correspondences, we adopt a proxy evaluation strategy using manually selected control points (as shown in Figure 11) as references. These points are selected by several EO experts to represent clearly identifiable and stable features across modalities. While this does not constitute absolute ground truth, the use of such human-verified correspondences provides a valuable reference, allowing us to assess how closely automated methods approximate human matching ability under multimodal and disaster conditions.
Table 12 presents the quantitative results of four UMIM methods on the Noto-Earthquake-2024 scene. These methods fall into two main categories: feature-based methods (LNIFT (Li et al., 2022) and SRIF (Li et al., 2023a)), which rely on sparse keypoint detection and matching, and area-based methods (FLSS (Ye et al., 2017) and HOPC (Ye et al., 2019)), which operate on local regions to estimate correspondences. A detailed description of each method is provided in Appendix H. Our experimental results reveal that traditional feature-based methods generally fail to achieve successful automatic registration. We attribute this limitation to the large spatial extent of the scenes and the drastic cross-modal differences, which make direct keypoint matching highly error-prone and unreliable. In contrast, area-based methods, which rely on the similarity in a local region rather than point correspondence in the global level, can partially mitigate registration errors and achieve relatively more stable performance. However, these methods still face a critical limitation: they typically perform local matching within a fixed search window. This strategy is inherently inadequate for handling large displacements, especially in SAR imagery over mountainous or uneven terrain, where terrain-induced distortions can cause substantial pixel shifts that far exceed the local search range. Thus, automatic methods may serve as a useful auxiliary tool for preliminary matching or candidate generation, but human expertise remains indispensable for ensuring precise and reliable alignment in operational scenarios. We hope that the Bright benchmark and its challenging real-world settings will inspire the development of new UMIM methods that are more robust to large scene variations, sensitive to terrain-induced distortions, and ultimately capable of reducing human efforts in future operational EO-based disaster response workflows.
@@ 5.1 Limitation of Bright
The performance of ML/DL models heavily depends on the quality of the training data. The Bright dataset is designed to enable ML/DL models trained on it to generate building damage maps of practical significance for future disaster events. It is anticipated that Bright will serve as a benchmark for numerous subsequent studies and practical disaster relief applications. However, the Bright dataset also has some limitations.
@@
    Registration error. Bright involves both optical and SAR imagery covering the same locations. SAR images, in particular, can be distorted and stretched in certain areas. Despite thorough preprocessing, including manual alignment and cross-checking by multiple experts in EO data processing, minor alignment errors may still persist, as suggested by Table 4.
@@
    Label quality. The building polygons in Bright were manually annotated by expert annotators. Although manual labeling generally ensures high accuracy, minor errors in polygon boundaries are inevitable due to the complexity of building shapes and the variability in image resolution. These inaccuracies may slightly affect the performance of the models trained on Bright. Then, the extent of damage to buildings is assessed by experts through visual interpretation of optical EO data. This process is susceptible to occasional misjudgments, contributing to label noise.
@@
    Sample and regional imbalance. Although Bright is richly geographically diverse, it has the problem of regional imbalance in the number of labels. Some of its events have more tiles and building numbers and thus are more dominant in training and evaluation, e.g., Turkey-Earthquake-2023 (1,114 tiles) v.s. Hawaii-Wildfire-2023 (65 tiles) in Table 2. Although we used an additional event-level evaluation method, the effect of events, which account for a large percentage of the sample, on the model during the training phase is still not negligible. It may affect the generalizability of the trained model in real disaster scenarios. Apart from that, all disaster events are located near the equator or in the northern hemisphere, with no events from the southern hemisphere represented. This spatial bias could potentially limit the applicability of trained models to regions with different building styles, land cover patterns, or SAR imaging characteristics prevalent in the southern part of the globe.
Overall, despite the limitations, it is the first time such a multimodal VHR dataset has been constructed that caters to multimodal EO research under large-scale and diverse disaster context.
@@ 5.2 Significance of Bright
Timely disaster response is often hindered by delays in both EO data acquisition and damage interpretation workflows (Ye et al., 2024). Traditional expert-driven building damage mapping is time-consuming and not scalable. While ML and DL methods offer automated alternatives, their effectiveness remains limited by the scope and quality of available training data. Most existing open-source datasets are optical-only, restricting the models’ operational applicability in adverse weather and low-light conditions. As the first globally distributed multimodal dataset, Bright encompasses both pre-event optical imagery and post-event SAR imagery. This unique combination overcomes the limitations of optical EO data by enabling models trained on Bright to monitor disaster-stricken areas regardless of weather conditions or daylight. Compared to existing building damage datasets, Bright offers several distinct characteristics: multimodal data, VHR imagery with sub-meter spatial resolution, coverage of five types of natural disasters, as well as two kinds of human-made disasters, rich geographic diversity, and open access to the community. Due to these features, Bright is anticipated to serve as a benchmark for many subsequent studies and practical disaster relief applications.
@@
Beyond building damage assessment, Bright can also be used to support several research directions within the EO and vision community. In this work, we have applied Bright to evaluate the performance of several UDA and SSL methods for cross-event transfer. Additionally, we have demonstrated its applicability to UMCD and UMIM, showcasing its versatility as a benchmark for multiple EO challenges under real-world constraints. Furthermore, the dataset provides a strong foundation for broader multimodal EO research. Its high-quality annotations and geographic diversity make it directly applicable to tasks such as building footprint extraction, land cover mapping, height estimation, and even EO-based visual question answering (VQA). Researchers can also repurpose or extend Bright to create task-specific benchmarks, enabling flexible experimentation across tasks. Bright is also well-positioned to support the development of remote sensing foundation models, large-scale, pre-trained models designed to generalize across sensors, tasks, and regions (Wang et al., 2023; Hong et al., 2024). Its rich combination of modalities, spatial detail, and contextual diversity provides the kind of data diversity required to build such general-purpose, transferable models. This contribution is significant as the field moves toward creating versatile, scalable AI models that can be applied across different types of EO data and disaster scenarios (Li et al., 2024)
@@
Looking ahead, Bright can be further enhanced by incorporating additional modalities, such as elevation data or multi-band SAR/optical inputs, to support terrain-aware analysis and improve semantic reasoning. Future versions may also help fill current geographic gaps, including southern hemisphere regions, ensuring more globally representative coverage. Ultimately, We envision that Bright, true to its name, will bring even a glimmer of brightness to people in disaster-stricken areas by enabling more prompt and effective disaster response and relief.

### v4

@@ 3.6.3 Why is cross-event transfer challenging?
To better understand why cross-event generalization is difficult, we explore two fundamental factors rooted in the nature of real-world disaster response.
@@
    Inconsistent damage signatures across events. Figure 10 presents violin plots of SAR backscatter values for intact, damaged, and destroyed buildings across multiple event pairs. These plots reveal two key observations. First, even within the same disaster category (the first row), the pixel intensity distributions for damaged and destroyed buildings differ significantly between events. This indicates that SAR-based damage signatures are inconsistent across locations, possibly due to differences in urban layout, building materials, or sensor incidence angles. Secondly, the distribution shift becomes more pronounced across different types of disasters (the second row). For example, the signatures of destroyed buildings in wildfires notably differ from those in volcano eruptions, which also differ from floods or hurricanes. These variations reflect fundamental differences in damage mechanisms: buildings burned in wildfires, submerged in floods, or collapsed in earthquakes leave very different patterns in SAR backscatter. Such distributional discrepancies make it extremely difficult for models to generalize from one event to another. A model trained on one disaster might learn class boundaries (e.g., between “Damaged” and “Destroyed”) that do not transfer well to another disaster, especially when the visual and physical properties of damage are fundamentally different.
    Figure 11: Comparison of models’ best performance (mIoU) on test events versus the best checkpoints selected on validation sets under the cross-event transfer setting. Each point represents the performance on a single test event under cross-event transfer. The farther a point lies from the diagonal line, the larger the gap between the model’s selected performance and its true upper bound.
@@
    Lack of target domain supervision for model selection. Another critical challenge in cross-event setup is the absence of labeled target domain samples for model selection. In typical domain adaptation benchmarks in computer vision, a validation set of target domain is available to tune hyperparameters or select the best checkpoint. However, no such validation data can be assumed in disaster response scenarios. Figure 11 illustrates the resulting issue: the gap between the actual best mIoU on the test event and the mIoU of the checkpoint selected using a source domain validation set for several representative models under both zero-shot and one-shot settings. In the zero-shot setting, this performance gap is substantial across most models. This indicates that relying on source domain validation leads to suboptimal model selection due to poor reflection of the target distribution. The one-shot setting helps reduce this gap by enabling limited target-aware selection. However, the issue is not entirely resolved. Even with a few labeled samples, model instability and domain shift still make selection challenging.
Together, these two findings highlight that cross-event transfer is hindered not only by semantic and statistical shifts across disaster types but also by operational constraints that prevent ideal training and tuning. Future studies should consider both aspects, i.e., developing models that are robust to distributional variance and designing selection strategies that do not depend on target supervision, such as self-validation, early stopping heuristics, or domain-agnostic performance proxies (Yang et al., 2023).
3.7 Evaluation of unsupervised multimodal change detection methods
Unsupervised multimodal change detection (UMCD) plays a crucial role in post-disaster assessment by enabling rapid analysis of affected areas without requiring labeled data or prior model training (Lv et al., 2022). Current studies are mainly conducted on toy datasets with limited geographic diversity and scene complexity. These datasets are for general land cover changes. Whether these methods work in real-world disaster occurrence/building damage scenarios is still unknown. Bright offers a new opportunity to evaluate UMCD methods to provide insights into their robustness and scalability in disaster response scenarios. Six representative UMCD methods (Sun et al., 2021; Chen et al., 2022b, 2023; Han et al., 2024; Sun et al., 2024a; Liu et al., 2025) are evaluated in this work. The experimental setup is described in Appendix H.
Table 13: Results of representative unsupervised multimodal change detection methods. KC is the acronym of kappa coefficient. The highest values are highlighted in purple, and the second-highest results are highlighted in teal. The accuracies on the UMCD benchmark dataset are the accuracies on the four datasets presented in Figure 13, obtained from their literature. Details of methods and benchmark datasets are presented in Appendix H. The random guessing baseline is included to indicate the performance floor under the UMCD setup. The “-” symbol indicates that the corresponding method did not report results on that dataset in their original publications.
@@ 3.7 Evaluation of unsupervised multimodal change detection methods
    Limitations in traditional UMCD datasets. The existing UMCD datasets consist of only a handful of image pairs, often depicting simple land cover changes, such as urban expansion and deforestation, with relatively low-resolution imagery and limited geographic diversity, as shown in Figure 13 in Appendix H. These datasets fail to capture the complexity and variability found in real-world disaster scenarios. In contrast, Bright provides thousands of VHR multimodal image pairs across different types of disasters, significantly increasing the diversity of test cases and making it a more challenging benchmark for UMCD model evaluation.
@@
    Interference from non-disaster changes. Unlike previous UMCD benchmarks, where general land cover changes are the primary detection target, changes such as vegetation growth, water body shifts, and urban development may introduce noise and interfere with damage detection in real-world disaster scenarios. While prior UMCD studies treat such changes as valid detection targets, Bright requires methods to differentiate true building structural damage from irrelevant land cover variations, posing a unique challenge to current approaches.
@@
    Problematic evaluation protocol. Current UMCD research follows a problematic evaluation protocol where models are trained, hyperparameters are tuned, and then validated on the same dataset. This approach can actually lead to overfitting, artificially making the model look very accurate but not generalizable to real-world scenarios. In a real-world disaster scenario, there are no labeled samples available for hyperparameter tuning, which makes the current practice unrealistic. Bright exposes this limitation, as models now need to be trained or tuned before being tested directly on new, unseen data, requiring truly generalizable and adaptive learning strategies.
The results have highlighted the challenges of applying existing UMCD models to real-world disaster scenarios. Bright reveals significant limitations in current methodologies and presents new opportunities for future UMCD research.
3.8 Evaluation of unsupervised multimodal image matching methods
Precise image alignment is a critical prerequisite for any multimodal EO application. Bright offers a unique opportunity to evaluate the performance of existing UMIM algorithms under realistic, large-scale disaster conditions. Due to the lack of real pixel-level ground truth correspondences, we adopt a proxy evaluation strategy using manually selected control points (as shown in Figure 12 in Appendix C) as references. These points were selected by EO experts to represent identifiable and stable features across modalities. While this does not constitute absolute ground truth, using such human-verified correspondences provides a valuable reference. This allows us to assess how closely automated methods approximate human matching ability under multimodal and disaster conditions.
@@ 3.8 Evaluation of unsupervised multimodal image matching methods
Table 14 presents the quantitative results of four UMIM methods on the Noto-Earthquake-2024 scene. These methods fall into two main categories: feature-based methods (LNIFT (Li et al., 2022) and SRIF (Li et al., 2023a)), which rely on sparse keypoint detection and matching, and area-based methods (FLSS (Ye et al., 2017) and HOPC (Ye et al., 2019)), which operate on local regions to estimate correspondences. A detailed description of each method is provided in Appendix I. The experimental results reveal that traditional feature-based methods fail to achieve successful automatic registration. We attribute this limitation to the large spatial extent of the scenes and the drastic cross-modal differences, which make direct keypoint matching highly error-prone and unreliable.
In contrast, area-based methods, which rely on the similarity in a local region rather than point correspondence at the global level, can partially mitigate registration errors and achieve relatively more stable performance. However, these methods still face a critical limitation: they typically perform local matching within a fixed search window. This strategy is inherently inadequate for handling large displacements, especially in SAR imagery over mountainous or uneven terrain, where terrain-induced distortions can cause substantial pixel shifts that far exceed the local search range. Thus, automatic methods may serve as a useful auxiliary tool for preliminary matching or candidate generation, but human expertise remains indispensable for ensuring precise and reliable alignment in operational scenarios. We hope that Bright and its challenging real-world scenarios will inspire the development of new UMIM methods that are more robust to large-scale scene variations, sensitive to terrain-induced distortions, and ultimately capable of reducing human efforts in future operational EO-based disaster response workflows.
@@ 4.1 Limitation of Bright
We begin this subsection by acknowledging that the composition of the Bright dataset is fundamentally shaped by practical constraints in data availability. While Bright represents a significant step forward in assembling a large-scale, multimodal, and globally distributed dataset for disaster response, it is important to recognize several inherent limitations. These limitations arise not only from the scarcity of open-access VHR SAR imagery, especially over disaster-affected regions, but also from the challenges of manual annotation and the uneven distribution of events. To provide a clearer picture for potential users, we summarize these constraints in four aspects below.
@@
    Registration error. Bright dataset consists of optical and SAR images covering the same locations. SAR images, in particular, can be distorted and stretched in certain areas. Despite thorough preprocessing, including manual alignment and cross-checking by multiple experts in EO data processing, minor alignment errors may persist, as Table 4 suggests.
@@
    Label quality. The building polygons in Bright were manually annotated by expert annotators. Although manual labeling generally ensures high accuracy, minor errors in polygon boundaries are inevitable due to the complexity of building shapes and the variability in image resolution. These inaccuracies may slightly affect the performance of the models trained on Bright. Furthermore, experts assessed the extent of damage to buildings through visual interpretation of optical EO data. This process is susceptible to occasional misjudgments, contributing to label noise.
@@
    Sample and regional imbalance. Although Bright is rich in geographically diverse data, it has the problem of regional imbalance in the number of labels. Some of its events have more tiles and building numbers, thus, are more dominant in training and evaluation, e.g., Turkey-Earthquake-2023 (1,114 tiles) v.s. Hawaii-Wildfire-2023 (65 tiles) in Table 2. To address this, we used an additional event-level evaluation method. However, the effect of events, which account for a large percentage of the sample, on the model during the training phase is still not negligible. This may affect the generalizability of the trained model in real-world disaster scenarios. In addition, all the disaster events in Bright are located near the equator or in the northern hemisphere, with no events from the southern hemisphere. This spatial bias could potentially limit the applicability of trained models to regions with different building styles, land cover patterns, or SAR imaging characteristics prevalent in the southern part of the globe.
  4. 4.
    Modality and temporal scope. The dataset’s scope is defined by two key characteristics of the available data. First, it exclusively utilizes single-polarization SAR imagery. The current version lacks the more informative multi-polarization or dense time-series SAR data, which, if available, could enable more nuanced damage characterization and long-term recovery monitoring, respectively. Second, the dataset’s temporal coverage is concentrated on events from 2020 onwards. This is a direct consequence of its reliance on modern commercial VHR SAR providers (Capella Space and Umbra), whose open-data initiatives largely commenced around that time.
Overall, despite these limitations, it is the first time such an open multimodal VHR dataset has been constructed for multimodal EO research with a large-scale and diverse disaster context.
@@ 4.2 Significance of Bright
Delays in both EO data acquisition and damage interpretation workflows often hinder timely disaster response (Ye et al., 2024). Traditional expert-driven building damage mapping is time-consuming and not scalable. While ML and DL methods offer automated alternatives, their effectiveness remains limited by the scope and quality of available training data. Most existing open-source datasets are optical, restricting the models’ operational applicability in adverse weather and low-light conditions. As the first globally distributed multimodal dataset, Bright encompasses pre-event optical images and post-event SAR images. This unique combination overcomes the limitations of optical EO data by enabling models trained on Bright to monitor disaster-stricken areas regardless of weather conditions or daylight. Compared to existing building damage datasets, Bright offers several distinct characteristics: multimodal data, VHR imagery with sub-meter spatial resolution, coverage of five types of natural disasters and two human-made disasters, rich geographic diversity, and open access to the community. Due to these features, Bright is anticipated to serve as a benchmark for many future studies and practical disaster relief applications.
@@
Beyond building damage assessment, Bright can also support several research directions within the EO and vision community. In this work, we have applied Bright to evaluate the performance of several UDA and SSL methods for cross-event transfer. We also demonstrated its applicability to UMCD and UMIM, showcasing its versatility as a benchmark for multiple EO challenges under real-world constraints. Furthermore, the dataset provides a strong foundation for broader multimodal EO research. Its high-quality annotations and geographic diversity directly apply to tasks, such as building footprint extraction, land cover mapping, height estimation, and EO-based visual question answering (VQA). Researchers can also repurpose or extend Bright to create task-specific benchmarks, enabling flexible experimentation across tasks. Bright is also well-positioned to support the development of EO-based foundation models, large-scale pre-trained models designed to generalize across sensors, tasks, and regions (Wang et al., 2023; Hong et al., 2024). Its rich combination of modalities, spatial detail, and contextual diversity provides the data diversity required to build such general-purpose foundation models. This contribution is significant as the field moves toward creating versatile, scalable AI models that can be applied across different types of EO data and disaster scenarios (Li et al., 2024).
@@
Looking ahead, Bright can be further enhanced by incorporating additional modalities. For example, the inclusion of fully polarimetric SAR data would enable more nuanced damage classification than current single polarization SAR data by characterizing the different scattering properties of building materials and debris. Meanwhile, LiDAR data would offer precise 3D information to directly quantify structural collapse and enable a truly terrain-aware analysis. Future versions may also help fill current geographic gaps, including southern hemisphere regions, to ensure more globally representative coverage. Ultimately, we envision that Bright, true to its name, will bring even a glimmer of brightness to people in disaster-stricken areas by enabling more prompt and effective disaster response and relief.
5 Data and Code Availability
The Bright dataset is available at https://doi.org/10.5281/zenodo.14619797 (Chen et al., 2025). The code for training and testing benchmark methods (including code related to IEEE GRSS DFC 2025) is accessible at https://github.com/ChenHongruixuan/BRIGHT. Models’ checkpoints can be downloaded at https://doi.org/10.5281/zenodo.15349461.